(69) Why is it impossible to find the instantaneous rate of change of f(x)= x^2-9/x-3 at x=3? THE SLASH IS THE FRACTION SYMBOL.


1
Expert's answer
2022-02-23T10:20:11-0500

Explanations & Calculations


  • It is because at "\small x =3" the graph has a vertical asymptote.
  • So the graph approaches infinity as it reaches "\small x \to 3".
  • The rate of change behaves the same way "\small""\small f'(x)_{x=3} = \frac{df(x)}{dx}_{x=3}\to -\infty"
  • The rate of change is given by

"\qquad\qquad
\begin{aligned}
\small f'(x) &=\small 2x-\frac{9}{(x-3)^2}
\end{aligned}"